Calculate the motor horsepower if a 2 × 2 in. plain milling cutter with 15 teeth is making a .250-in.-deep cut, 2 in. wide at 120 rpm, with a feed rate of .00055 in./tooth, unit horsepower of 1, and 75% efficiency.
Answer #1

Feed rate per revolution = 0.00055*15 = 0.00825 in/rev

Feed rate per min = 0.00825*N = 0.00825*120 = 0.99 in/min

Material removal rate = MRR = feed rate/min * Area of cut = 0.99*0.250*2 = 0.495 in3/ min

Spindle power = MRR* unit power = 0.495 * 1 = 0.495 hp

Motor power = spindle power/ efficiency = 0.495/0.75 = 0.66 hp

Motor power = Pm = 0.66 hp : ANSWER